OpenStudy (anonymous):

LAST QUESTION What is the slope-intercept form of the function that contains the point (1, –2) and has a slope of –3?

OpenStudy (bibby):

start by plugging the information into the point-slope form equation \(y-y_1=m(x-x_!)\) where \((x_1,y_2)\) is the point the line goes through and m is the slope of the line, then solve for y

OpenStudy (bibby):

x1,y1*

OpenStudy (anonymous):

idk

OpenStudy (bibby):

what does that mean? read what I said. \(y-\color{green}{y_1}=\color{blue}m(x-\color{red}{x_1})\) \((\color{red}{1}, \color{green}{–2}) \) \( slope = \color{blue}{–3}\)

OpenStudy (bibby):

refresh to see the negatives

OpenStudy (anonymous):

y-(-2)=-3(x-1)

OpenStudy (anonymous):

it needs to be in y=mx+b form

OpenStudy (bibby):

correct. so simplify the left side. y-(-2)=y+2 then distribute the -3

OpenStudy (bibby):

that's why I said solve for y

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

y+2= -3x - -3

OpenStudy (bibby):

yep

OpenStudy (anonymous):

ok now what

OpenStudy (bibby):

solve for y? that means isolate y. get it a lone

OpenStudy (anonymous):

how do i do that

OpenStudy (bibby):

the same way you'd solve y+2=45

OpenStudy (anonymous):

i dont know how to isolate y

OpenStudy (bibby):

how would you do it in the "y+2=45"

OpenStudy (anonymous):

divide 2 by 45?

OpenStudy (anonymous):

can i add the 2 to b?

OpenStudy (bibby):

subtract 2. addition and subtraction are inverse operations

OpenStudy (bibby):

if you add something and then subtract the same thing, you end up with the same thing

OpenStudy (bibby):

x+4-4=x

OpenStudy (anonymous):

so y= -3x + 1 ?

OpenStudy (bibby):

y-(-2)=-3(x-1) y+2=-3x--3 y+2=-3x+3 y=-3x+1 yeah, good job

OpenStudy (bibby):

another way of doing this is using y=mx+b where m is the slope and b is the y intercept

OpenStudy (bibby):

the y intercept is where x=0, so you'd have to find (0,y) first since the slope is -3, if we go back one x unit, we'd subtract -3 from the y (1, –2) (0, 1) and so we have y=-3x+1

OpenStudy (bibby):

god dammit I hate this site